Determine whether the function T: R

3 →R

3 by 𝑇 ([

𝑥

𝑦

𝑧

]) = [

𝑥 + 𝑦 − 𝑧

2𝑥𝑦

𝑥 + 𝑧 + 1

] is a linear 

transformation between vector space.



1
Expert's answer
2021-05-19T17:17:35-0400

The given function-


"T(\begin{bmatrix} x\\y\\z

 \end{bmatrix})=\begin{bmatrix}x+y-z\\2xy\\x+z+1\end{bmatrix}"


    Let"\alpha=\begin{bmatrix} \alpha_1\\\alpha_2\\\alpha_3

 \end{bmatrix},\beta=\begin{bmatrix} \beta_1\\\beta_2\\\beta_3

 \end{bmatrix}\in R^3"


"T(\alpha+\beta)=T(\begin{bmatrix} \alpha_1\\\alpha_2\\\alpha_3

 \end{bmatrix}+\begin{bmatrix} \beta_1\\\beta_2\\\beta_3

 \end{bmatrix})

\\[9pt]

 =T(\begin{bmatrix} \alpha_1+\beta_1\\\alpha_2+\beta_2\\\alpha_3+\beta_3

 \end{bmatrix})

\\[9pt]

=\begin{bmatrix} \alpha_1+\beta_1+\alpha_2+\beta_2-(\alpha_3+\beta_3)\\2(\alpha_1+\beta_1)(\alpha_2+\beta_2)\\\alpha_1+\beta_1+\alpha_3+\beta_3+1

 \end{bmatrix}\\[9pt]



=\begin{bmatrix} \alpha_1+\alpha_2-\alpha_3\\2\alpha_1 \alpha_2\\\alpha_1+\alpha_3+1

 \end{bmatrix}+\begin{bmatrix} \beta_1+\beta_2-\beta_3\\2\beta_1 \beta_2+2(\alpha_1 \beta_2+\beta_1 \alpha_2)\\\beta_1+\beta_3

 \end{bmatrix}\\[9pt]\neq T(\alpha)+T(\beta)"




Since, "T(\beta)=\begin{bmatrix} \alpha_1+\alpha_2-\alpha_3\\2\alpha_1 alpha_2\\\alpha_1+alpha_3+1

 \end{bmatrix}\neq\begin{bmatrix} \beta_1+\beta_2-\beta_3\\2\beta_1 \beta_2+2(\alpha_1 \beta_2+\beta_1 \alpha_2)\\\beta_1+\beta_3

 \end{bmatrix}"


Therefore T is not liner transformation between vector space.